Consider a Voltaic so made up of copper and sodium half cells. Write the half reactions, the balanced overall reactions, and determine the cell potential under standard conditions is this reaction spontaneous or nonspontaneous
1
Expert's answer
2020-04-19T07:22:04-0400

Half-reactions:

cathode: Cu2+ + 2 e- = Cu

anode: Na - e- = Na+

Overall:

Cu2+ + 2 Na = Cu + 2 Na+ (ionic form)

CuCl2 + 2 Na = Cu + 2 NaCl (example of complete form)

Cell standard potential:

E0 = E0(Cu2+/Cu) - E0(Na+/Na) = 0.337 + 2.710 = 3.047 V

Reaction is spontaneous in the forward direction (because E0 > 0).